find the coordinates of the vertex and the equation of the axis of symmerty for the parabola with the equation 2x^2 + 2x - y = -3
Found 3 solutions by stanbon, Alan3354, RAY100:
Answer by stanbon(75887)

find the coordinates of the vertex and the equation of the axis of symmerty for the parabola with the equation 2x^2 + 2x - y = -3
----------------------
Rearrange:
2x^2 + 2x + ? = y-3
---
2(x^2 + x + (1/2)^2) = y-3+2(1/2)^2
---
2(x+(1/2))^2 = y-(11/4)
----
Vertex=(-1/2 , 11/4)
Axis of symmetry: x = -1/2
=================================
Cheers,
Stan H.

Answer by Alan3354(69443)

find the coordinates of the vertex and the equation of the axis of symmerty for the parabola with the equation 2x^2 + 2x - y = -3
y = 2x^2 + 2x + 3
-----------------
The axis of symmetry is x = -b/2a = -2/4
x = -1/2 is the axis of symmetry.
-------------
At x = -1/2: y = 1/2 - 1 + 3
y = 5/2
Vertex at (-1/2,5/2)

Answer by RAY100(1637)

2x^2 +2x =y-3
.
2(x^2 +x ) =y-3,,,,factor out 2
.
2(x^2 +x +1/4) =y-3+2(1/4),,,,,complete square and add compensating to rt side
.
2(x+1/2)^2 = y-2.5
.
Remember the std form of a parabola is (y-k) = A(x-h)^2,,,,with(h,k) as vertex
.
Vertex in this case is,,,,(-1/2, 2.5)
.
line of symmetry is at ,,,,x=-1/2
.
A is positive ,,,(+2),,,therefore y parabola pointing up
.